# The inclusive Synthetic Control Method

### v1

@@ 3 An example with a single “potentially affected” unit
where ($\widehat{\theta}_{t}$) is the estimated effect for the “main treated”; ($\widehat{\gamma}_{t}$) is the estimated effect for the “potentially affected” unit; ($\widehat{w}_{2}$) and ($\widehat{l}_{1}$) are the estimated weights; ($\theta_{t}$) and ($\gamma_{t}$) are the unknown effects.
@@
In this case, it is easy to see that $\det(\widehat{\Omega})$ is always different from zero, except when $\widehat{w}_{2}$ = $\widehat{l}_{1}=1$. Thus, our effects of interest are always identified unless the “main treated” gives weight 1 to the other “potentially affected” unit, which in turn gives weight 1 to the “main treated”. This would be the case, for example, if there are no “pure control” units.
An interesting special case is when we do not need to include the “main treated” in the donor pool of the “potentially affected” unit. In this case, the system of equations further simplifies to
@@ 4 Bias comparison between iSCM and SCM for $\bm{m=1}$
The bias of $\widehat{\theta}_{t}^{iSCM}$ approaches zero as the estimation biases $B_{sc}^{1}$ and $B_{sc}^{2}$ approach zero. In contrast, if we include “potentially affected” units, $Bias\left(\widehat{\theta}_{t}\right)=-\widehat{w}_{2}\gamma_{t}$, even when $B_{sc}^{1}$ approaches zero.
@@
In case it is not necessary to include the “main treated” in the donor pool of the “potentially affected” unit, we have
@@
5 iSCM vs “restricted” SCM
When using SC-type estimators, it is advisable to include in the donor pool units with similar characteristics and those possibly affected by similar shocks as the treated unit. Often, these units are either directly or indirectly affected by the intervention. For instance, it is likely that other treated units or units “potentially affected” by spillover effects are the closest (geographically and/or economically) to the “main treated” unit. For example, Abadie (2021) proposes including units “potentially affected” by spillover in the donor pool, even if they induce a bias.
Indeed, discarding “potentially affected” units could substantially decrease the match quality between the characteristics of the treated and the synthetic control, as those units are typically the closest to the treated. Therefore, choosing between our iSCM and a “restricted” version of SCM, where “potentially affected” units are excluded, depends on the importance of the excluded units in creating the synthetic control.
In the Appendix’s Section A.2, we demonstrate that iSCM may be superior to a “restricted” SCM even in scenarios where both “main treated” and “potentially affected” units are within the “pure control” units’ convex hull. In this scenario, a “restricted” SCM will potentially work, however, if the “potentially affected” units are the closest to the “main treated”, excluding them might increase interpolation bias. Additionally, we explore cases more common in empirical settings, where either or both the “main treated” and “potentially affected” units fall outside the convex hull.
To understand whether we expect a large bias by excluding “potentially affected” units, before implementing our iSCM, we advise taking the following steps to assess the potential gains of using our method instead of running a “restricted” SCM. When the SC-type estimator used is the one proposed by Abadie et al. (2010), we suggest following the procedure described below before adopting our method.
@@
    Implement SCM, including units “potentially affected” by the intervention (i.e., other treated units or units affected by spillovers).
  • If all “potentially affected” units receive low or zero weights, they induce a negligible bias and can be used as “pure controls” in a SCM.
  • If some of the “potentially affected” units receive high weights, proceed with step 2.
@@
    Implement the “restricted” SCM, i.e., excluding units “potentially affected” by the intervention and:
    1. (a)
@@
  • If ($X_{1}-X_{0}^{res}\widehat{W}^{res})\approx(X_{1}-X_{0}\widehat{W}^{unres}$) and $RMSPE^{res}\approx RMSPE^{unres}$, we advise using the “restricted” SCM as main specification and iSCM as a robustness check [^6] Alternatively, one could divide the pre-intervention period into a training period and a validation period, estimate the two models during the training period, and compare their RMSPEs in the validation period. Note that the “restricted” SCM does not have by construction an RMSPE that is greater than or equal to that of the “unrestricted” one, see for example Section 7..
  • If ($X_{1}-X_{0}^{res}\widehat{W}^{res})>(X_{1}-X_{0}\widehat{W}^{unres}$) and/or $RMSPE^{res}>RMSPE^{unres}$, we advise using our iSCM.
Repeat these steps for each “potentially affected” unit as if it were the “main treated”. A similar procedure can be used for other SC-type estimators. In case there is no substantial gain from including the treated in the donor pool of “potentially affected” units, our iSCM becomes easier to implement, as shown in Equation 7. Finally, it’s worth comparing the results of the three methods, keeping in mind that the “unrestricted” SCM is biased if any of the units in the donor pool receiving non-negligible weight is affected by the treatment.
@@ 6 Inference
Dealing with only a small number of units makes inference for synthetic control-based methods like ours complicated. We can, however, easily adapt existing methods to our setting. The most popular choice is to implement permutation tests. Abadie et al. (2010) and Abadie et al. (2015) propose placebo tests in time, i.e., reassigning the intervention artificially before its real implementation, and placebo tests in-space, i.e., reassigning the intervention artificially for units in the control group. The latter approach is often preferred because of possible shocks that might have occurred in the past that affect units differently. Placebo tests in space measure the statistical significance of the effect through the ratio between the RMSPE in the post-treatment period and in the pre-treatment period.
To run the in-space placebo test for $\widehat{\theta}_{t}^{iSCM}$, all we need to do is subtract the estimated effects from the outcomes of “potentially affected” units. That is, we replace $Y_{it}$ with $Y_{it}-\widehat{\gamma}_{it}^{iSCM}$, for all $i=2,\ldots,m$, and $t>T_{0}$. Similarly, to construct a placebo test for $\widehat{\gamma}_{it}^{iSCM}$ for a generic “potentially affected” unit $i$, we need to subtract the estimated effects from the outcomes of all other “potentially affected” units as well as the “main” treated unit, leaving the outcome of unit $i$ untouched.
This idea can easily be applied to other inference procedures available in the literature (see, e.g., Cao and Dowd 2019; Chernozhukov et al. 2021; Firpo and Possebom 2018; Gobillon and Magnac 2016; Li 2019). For example, the inference procedure of Chernozhukov et al. (2021) can be easily adapted to our framework by taking into account that the post-intervention observed outcome of each “potentially affected” unit includes its treatment effect. This has to be considered both in the first step, i.e., in the construction of the data under the null hypothesis and in the second step to compute the SCM residuals.
@@ 7 Empirical example
As discussed in Abadie et al. (2015) and Abadie and L’Hour (2021), German reunification could have had negative spillover effects on Austria’s economic growth because West Germany diverted demand and investment from Austria to East Germany. Austria has historically had tight links with Germany: the two countries share the same language and, to a great extent, a common history. In 1938, Austria was annexed by the Third Reich, which benefited from Austria’s raw materials and labor to complete German rearmament. In 1945, Austria was separated from Germany. However, the economic cooperation between Austria and West Germany continued during the Cold War. Given these strong cultural and economic ties, it is arguably important to include Austria in the donor pool when constructing a synthetic version of West Germany. Therefore, our iSCM is very well suited for estimating the impact of the German reunification not only on West Germany but also on Austria. [^7] Given that other European countries receive very little weight, the impact of potential spillover effects on those countries would likely be negligible, as shown in Lemma 1. Results for the case in which Switzerland and the Netherlands are considered as “potentially affected” units are available from the authors upon request.
@@
We use the same country-level panel data of Abadie et al. (2015). The data cover the period 1960-2003, with the post-intervention period starting in 1990. In addition to Austria, the remaining “pure control” countries in the donor pool include 15 other OECD countries: Australia, Belgium, Denmark, France, Greece, Italy, Japan, the Netherlands, New Zealand, Norway, Portugal, Spain, Switzerland, the United Kingdom, and the United States. The outcome variable is real GDP at Purchasing Power Parity (PPP) per capita measured in 2002 USD.
We start by implementing the procedure described in Section 5. First, we replicate the SCM estimate of Abadie et al. (2015) and check the weight assigned to Austria [^8] We refer to Abadie et al. (2015) for a detailed discussion of the estimation procedure.. Given that Austria does receive the highest weight (42%) we can proceed to step 2, i.e., implementing the “restricted” SCM, i.e., excluding Austria from the donor pool, keeping the same specification and estimation procedure used for the “unrestricted” SCM. Table 1 suggests that the “unrestricted” synthetic version of West Germany (second column) is much closer in therms of observable characteristics to West Germany (first column) than the “restricted” version (third column).
@@
Next, we compare the pre-intervention RMSPEs of the “unrestricted” and “restricted” SCM. The RMSPE of the latter (270.74) is larger than the one of the former (119.07). Therefore, we expect iSCM to perform better than “restricted” SCM. We now repeat the same procedure to decide whether West Germany should be included in synthetic Austria’s donor pool. First, we check whether West Germany receives a non-negligible weight. To construct synthetic Austria, we must use a slightly different specification than the one used for creating synthetic West Germany. In particular we are not able to chose the weights assigned to the predictors using the sample splitting methods described in Abadie et al. (2015). As described in Gehler and Graf (2018), in 1980, not long before the sample split cut-off, Austria provided several loans to East Germany, and in return, Austrian nationalized industries received large-scale orders. This most likely stimulated Austria’s exports and contributed to job creation in its industries. Thus, the sample split procedure might catch the effect of this economic shock. This is corroborated by the fact that using this method to chose the predictor weights leads to a poor pre-intervention fit. For this reason, we decided to follow the data driven procedure suggested by Abadie et al. (2010) instead.
@@
As shown in Table 2, Synthetic Austria gives the highest weight to West Germany. Thus, we can proceed to step 2, i.e., checking how well “restricted” (excluding West Germany) and “unrestricted” synthetic Astria matches real Austria’s observable characteristics and comparing the pre-intervention RMPEs of the two specifications.
@@
Table 1 suggests that the “unrestricted” synthetic Austria (second column) does a better job of reproducing Austria’s (first column) pre-reunification predictors than the “restricted” version (third column), except for GDP per capita. The pre-intervention RMSPE of the “restricted” SCM (181.22) is slightly lower than the one of the “unrestricted” version (194.67). However, the difference is rather small. Taking everything into account, we argue that it is better to include West Germany in Austria’s donor pool.
@@
Now we can use the “unrestricted” SCM estimates of $\widehat{\theta_{t}}$, $\widehat{\gamma_{t}}$, the weight assigned to Austria $\hat{w}_{A}=0.42$, and the weight assigned to West Germany $\hat{l}_{WG}=0.33$ to estimate $\theta_{t}$, and $\gamma_{t}$ as
@@
Figure 1 shows our main results. In Panel 1a and 1b we report the gaps estimated by the “unrestricted” SCM, the “restricted” SCM, and our iSCM, for West Germany and Austria, respectively [^9] Notice that the “unrestricted” SCM and the iSCM are identical in the pre-treatment period because the spillover effect due to German reunification only happens after the intervention.. In Panel 1c we plot the GDP per capita trajectories (1960–2003) of Real West Germany, its “unrestricted” synthetic version, its “restricted” synthetic version, and its inclusive synthetic version. Panel 1d produces an analogous plot for Austria.
@@
If we look at the estimated spillover effects for Austria we draw similar conclusions. In the post-reunification period, iSCM estimates a decrease in per capita GDP of only up to 708 USD. However, this effect is unlikely to be statistically significant (see Figure Panel ). It is important to notice that both the “unrestricted” SCM and the “restricted” SCM estimate a positive spillover effect for Austria, up to 894 and 1350 USD, respectively. However, we know a priori that the “unrestricted” SCM is biased and its bias is $-\hat{l}_{WG}\theta_{t}$. Given we have strong evidence that $\theta_{t}$ is negative, we do expect the “unrestricted” SCM to overestimate $\gamma_{t}$. Thus, the fact that both the “unrestricted” and the “restricted” SCM estimate positive spillover effects underscores the necessity of including West Germany in the donor pool and employing our iSCM in this application.
@@
Finally, Panels 2a and 2b show the ratios between the RMSPEs in the post- and pre-reunification of West Germany’s donor pool and Austria’s donor pool, respectively. We can observe that West Germany’s value is very high and is the largest compared to any other countries in the donor pool. On the other hand, Austria’s RMSPE ratio is the second lowest indicating the spillover effect is not significant, as we mentioned above.
(a) RMSPE ratios West Germany
(b) RMSPE ratios Austria

### v2

@@ 3 An example with a single ”potentially affected” unit
where ($\widehat{\theta}_{t}$) is the estimated effect for the ”main treated”; ($\widehat{\gamma}_{t}$) is the estimated effect for the ”potentially affected” unit; ($\widehat{w}_{2}$) and ($\widehat{l}_{1}$) are the estimated weights; ($\theta_{t}$) and ($\gamma_{t}$) are the unknown effects.
@@
In this case, it is easy to see that $\det(\widehat{\Omega})$ is always different from zero, except when $\widehat{w}_{2}$ = $\widehat{l}_{1}=1$. Thus, our effects of interest are always identified unless the ”main treated” gives weight 1 to the other ”potentially affected” unit, which in turn gives weight 1 to the ”main treated”. This would be the case, for example, if there are no ”pure control” units.
An interesting special case is when we do not need to include the ”main treated” in the donor pool of the ”potentially affected” unit. In this case, the system of equations further simplifies to
@@ 4 Bias comparison between iSCM and SCM for $\bm{m=1}$
The bias of $\widehat{\theta}_{t}^{iSCM}$ approaches zero as the estimation biases $B_{sc}^{1}$ and $B_{sc}^{2}$ approach zero. In contrast, if we include ”potentially affected” units, $Bias\left(\widehat{\theta}_{t}\right)=-\widehat{w}_{2}\gamma_{t}$, even when $B_{sc}^{1}$ approaches zero.
@@
In case it is not necessary to include the ”main treated” in the donor pool of the ”potentially affected” unit, we have
@@
5 iSCM vs ”restricted” SCM
When using SC-type estimators, it is advisable to include in the donor pool units with similar characteristics and those possibly affected by similar shocks as the treated unit. Often, these units are either directly or indirectly affected by the intervention. For instance, it is likely that other treated units or units ”potentially affected” by spillover effects are the closest (geographically and/or economically) to the ”main treated” unit. For example, Abadie (2021) proposes including units ”potentially affected” by spillover in the donor pool, even if they induce a bias, which our iSCM eliminates.  [^4] In the appendix (Section A.3), we show that iSCM may still be superior to ”restricted” SCM even in scenarios where both ”main treated” and ”potentially affected” units are within the ”pure control” units’ convex hull. In this case, a ”restricted” SCM will potentially work, however, if the “potentially affected” units are the closest to the “main treated”, excluding them might increase interpolation bias. Additionally, we explore cases more common in empirical settings, where either or both the ”main treated” and ”potentially affected” units fall outside the convex hull. For example, if $m=1$, as we have shown in Section 4, iSCM bias is $\frac{-B_{sc}^{1}-\widehat{w}_{2}B_{sc}^{2}}{1-\widehat{w}_{2}\widehat{l}_{1}}$. Let the approximation error of $Y_{1t}^{N}$ obtained with a the ”restricted” SCM be $B_{rSC}$. Then the ”restricted” SCM bias will be $-B_{rSC}$. It is hard a priori to judge which method has a larger bias as this depends on the sign of $-B_{rSC}-\frac{-B_{sc}^{1}-\widehat{w}_{2}B_{sc}^{2}}{1-\widehat{w}_{2}\widehat{l}_{1}}$ which is equal to the sign of $(B_{sc}^{1}+\widehat{w}_{2}B_{sc}^{2})-(1-\widehat{w}_{2}\widehat{l}_{1})B_{rSC}$. Thus the difference in the biases of iSCM and ”restricted” SCM is a function of three unknown approximation biases.
Therefore, we suggest implementing our iSCM and comparing it to the ”restricted” SCM, i.e., excluding potentially affected units. To determine which of the two methods should be used as the main specification, we advise taking the following steps.  [^5] We recommend implementing and reporting the results of both methods to avoid potential pretest bias (see Roth 2022).
@@
    Check whether the ”potentially affected” units receive substantial weights. If they do not, both methods should perform similarly.
@@
If ($X_{1}-X_{0}^{res}\widehat{W}^{res})\approx(X_{1}-X_{0}\widehat{W}^{unres}$) and $RMSPE^{res}\approx RMSPE^{unres}$, then the “restricted” SCM is preferable  [^6] Alternatively, one could divide the pre-intervention period into a training period and a validation period, estimate the two models during the training period, and compare their RMSPEs in the validation period. Note that the ”restricted” SCM does not have by construction an RMSPE that is greater than or equal to that of the ”unrestricted” one, see for example Section 7.. If ($X_{1}-X_{0}^{res}\widehat{W}^{res})>(X_{1}-X_{0}\widehat{W}^{unres}$) and/or $RMSPE^{res}>RMSPE^{unres}$, we advise using our iSCM as main specification. Note that, unlike a regression model, the ”restricted” SCM might still achieve a lower RMSPE than the ”unrestricted” one (see, for example, our empirical application), making this comparison meaningful.
Repeat these steps for each ”potentially affected” unit as if it were the ”main treated”. In case there is no substantial gain from including the treated in the donor pool of ”potentially affected” units, our iSCM becomes easier to implement, as shown in Equation 7. Finally, it is worth comparing the results of iSCM and the ”restricted” SCM with those of the ”unrestricted” SCM, bearing in mind that the latter can be biased if any of the units in the donor pool receiving non-negligible weight are affected by the treatment.
@@ 6 Inference
Dealing with only a small number of units makes inference for synthetic control-based methods like ours complicated. We can, however, easily adapt existing methods to our setting. The most popular choice is to implement permutation tests. Abadie et al. (2010) and Abadie et al. (2015) propose placebo tests in time, i.e., reassigning the intervention artificially before its real implementation, and placebo tests in-space, i.e., reassigning the intervention artificially for units in the control group. Placebo tests in space measure the statistical significance of the effect through the ratio between the RMSPE in the post-treatment period and in the pre-treatment period.
To run the in-space placebo test for $\widehat{\theta}_{t}^{iSCM}$, we simply need to subtract the estimated effects from the outcomes of ”potentially affected” units. Specifically, we replace $Y_{it}$ with $Y_{it}-\widehat{\gamma}_{it}^{iSCM}$ for all $i=2,\ldots,m$, and $t>T_{0}$, as follows:
$r_{1}=\frac{\left(\frac{1}{T-T_{0}}\sum_{t=T_{0}+1}^{T}\left(Y_{1t}-\left(\sum_{j=m+1}^{J}\widehat{w}_{j}Y_{jt}+\sum_{j=2}^{m}\widehat{w}_{j}\left(Y_{jt}-\widehat{\gamma}_{jt}^{iSCM}\right)\right)\right)^{2}\right)^{1/2}}{\left(\frac{1}{T_{0}}\sum_{t=1}^{T_{0}}\left(Y_{1t}-\sum_{j=2}^{J}\widehat{w}_{j}Y_{jt}\right)^{2}\right)^{1/2}},$
Similarly, to construct a placebo test for $\widehat{\gamma}_{it}^{iSCM}$ for a generic ”potentially affected” unit $i$, we need to subtract the estimated effects from the outcomes of all other ”potentially affected” units as well as the ”main” treated unit, leaving the outcome of unit $i$ untouched. For $i=2,\ldots,m$, we have:
$r_{i}=\frac{\left(\frac{1}{T-T_{0}}\sum_{t=T_{0}+1}^{T}\left(Y_{it}-\left(\sum_{j=m+1}^{J}\widehat{l}^{i}_{j}Y_{jt}+\sum_{j\in M\backslash\{{i}\}}\widehat{l}^{i}_{j}\left(Y_{jt}-\widehat{\gamma}_{jt}^{iSCM}\right)+\left(\widehat{l}^{i}_{1}Y_{1t}-\widehat{\theta}^{iSCM}_{1t}\right)\right)\right)^{2}\right)^{1/2}}{\left(\frac{1}{T_{0}}\sum_{t=1}^{T_{0}}\left(Y_{it}-\sum_{j\neq i}\widehat{w}_{j}Y_{jt}\right)^{2}\right)^{1/2}}.$
This idea can easily be applied to other inference procedures available in the literature (see, e.g., Cao and Dowd 2019; Chernozhukov et al. 2021; Firpo and Possebom 2018; Gobillon and Magnac 2016; Li 2019). For example, the inference procedure of Chernozhukov et al. (2021) can be easily adapted to our framework by taking into account that the post-intervention observed outcome of each ”potentially affected” unit includes its treatment effect. This has to be considered both in the first step, i.e., in the construction of the data under the null hypothesis and in the second step to compute the SCM residuals.
@@ 7 Empirical example
As discussed in Abadie et al. (2015) and Abadie and L’Hour (2021), German reunification could have had negative spillover effects on Austria’s economic growth because West Germany diverted demand and investment from Austria to East Germany. Austria has historically had tight links with Germany: the two countries share the same language and, to a great extent, a common history. In 1938, Austria was annexed by the Third Reich, which benefited from Austria’s raw materials and labor to complete German rearmament. In 1945, Austria was separated from Germany. However, the economic cooperation between Austria and West Germany continued during the Cold War. Given these strong cultural and economic ties, it is arguably important to include Austria in the donor pool when constructing a synthetic version of West Germany. Therefore, our iSCM is very well suited for estimating the impact of the German reunification not only on West Germany but also on Austria.
@@
We use the same country-level panel data of Abadie et al. (2015). The data cover the period 1960-2003, with the post-intervention period starting in 1990. In addition to Austria, the remaining ”pure control” countries in the donor pool include 15 other OECD countries: Australia, Belgium, Denmark, France, Greece, Italy, Japan, the Netherlands, New Zealand, Norway, Portugal, Spain, Switzerland, the United Kingdom, and the United States. The outcome variable is real GDP at Purchasing Power Parity (PPP) per capita measured in 2002 USD.
@@
We replicate the SCM estimate of Abadie et al. (2015) and check the weight assigned to Austria. [^7] We refer to Abadie et al. (2015) for a detailed discussion of the estimation procedure. The fact that Austria is potentially affected by spillover effects implies that including it in the donor pool might induce bias. [^8] Given that other European countries receive very little weight, the impact of potential spillover effects on those countries would likely be negligible, as shown in Lemma 1. Results for the case in which Switzerland and the Netherlands are considered as ”potentially affected” units are available from the authors upon request. As we suggest in Section 5, to decide which of the two methods is preferable, we need to implement both our iSCM and the ”restricted” SCM, i.e., excluding Austria from the donor pool, keeping the same specification and estimation procedure used for the ”unrestricted” SCM. Austria receives the highest weight (42%) thus iSCM and ‘unrestricted” SCM might give different results. Table 1 suggests that the “unrestricted” synthetic version of West Germany (second column) is much closer in terms of observable characteristics to West Germany (first column) than the “restricted” version (third column). Next, we compare the pre-intervention RMSPEs of the “unrestricted” and “restricted” SCM. The RMSPE of the latter (270.74) is larger than the one of the former (119.07). Therefore, we expect iSCM to perform better than “restricted” SCM. We now repeat the same procedure to decide whether West Germany should be included in synthetic Austria’s donor pool. First, we check whether West Germany receives a non-negligible weight. To construct synthetic Austria, we must use a slightly different specification than the one used for creating synthetic West Germany. In particular, we are not able to choose the weights assigned to the predictors using the sample splitting methods described in Abadie et al. (2015). As described in Gehler and Graf (2018), in 1980, not long before the sample split cut-off, Austria provided several loans to East Germany, and in return, Austrian nationalized industries received large-scale orders. This most likely stimulated Austria’s exports and contributed to job creation in its industries. Thus, the sample split procedure might catch the effect of this economic shock. This is corroborated by the fact that using this method to choose the predictor weights leads to a poor pre-intervention fit. For this reason, we decided to follow the data driven procedure suggested by Abadie et al. (2010) instead.
@@
As shown in Table 2, Synthetic Austria gives the highest weight to West Germany (33%). Thus, we can proceed to checking how well ”restricted” (excluding West Germany) and ”unrestricted” synthetic Austria matches real Austria’s observable characteristics and comparing the pre-intervention RMPEs of the two specifications.
@@
Table 1 suggests that the “unrestricted” synthetic Austria (second column) does a better job of reproducing Austria’s (first column) pre-reunification predictors than the “restricted” version (third column), except for GDP per capita. The pre-intervention RMSPE of the “restricted” SCM (181.22) is slightly lower than the one of the “unrestricted” version (194.67). However, the difference is rather small. Taking everything into account, we argue that iSCM is preferable to the “restricted” SCM.
@@
Now we can use the ”unrestricted” SCM estimates of $\widehat{\theta_{t}}$, $\widehat{\gamma_{t}}$, the weight assigned to Austria $\hat{w}_{A}=0.42$, and the weight assigned to West Germany $\hat{l}_{WG}=0.33$ to estimate $\theta_{t}$, and $\gamma_{t}$ as
@@
Figure 1 shows our main results. In Panel 1a and 1b we report the gaps estimated by the ”unrestricted” SCM, the ”restricted” SCM, and our iSCM, for West Germany and Austria, respectively [^9] Notice that the “unrestricted” SCM and the iSCM are identical in the pre-treatment period because the spillover effect due to German reunification only happens after the intervention.. In Panel 1c we plot the GDP per capita trajectories (1960–2003) of Real West Germany, its ”unrestricted” synthetic version, its “restricted” synthetic version, and its inclusive synthetic version. Panel 1d produces an analogous plot for Austria.
@@
If we look at the estimated spillover effects for Austria we draw similar conclusions. In the post-reunification period, iSCM estimates a decrease in per capita GDP of only up to 708 USD. However, this effect is unlikely to be statistically significant (see Figure 3 Panel 3b in the appendix). It is important to notice that both the ”unrestricted” SCM and the ”restricted” SCM estimate a positive spillover effect for Austria, up to 894 and 1350 USD, respectively. However, we know a priori that the ”unrestricted” SCM is biased and its bias is $-\hat{l}_{WG}\theta_{t}$. Given we have strong evidence that $\theta_{t}$ is negative, we do expect the ”unrestricted” SCM to overestimate $\gamma_{t}$. Thus, the fact that both the ”unrestricted” and the ”restricted” SCM estimate positive spillover effects underscores the necessity of including West Germany in the donor pool and employing our iSCM in this application.
@@
7.1 Applying the inclusive Penalized SCM
In this section, we re-estimate the effects on West Germany and Austria using the Penalized Synthetic Control Method.
We adopt the same specification used in the original SCM analysis, selecting the penalty parameter, $\lambda$, through cross-validation. We retain the same weight matrix $V$. Although one might expect that a method improving pre-intervention fit, such as penalized SCM, would mitigate the impact of excluding Austria from the donor pool, the results suggest otherwise. Figure 2 shows that the pre-treatment fit is substantially worse in the restricted penalized SCM version for both countries, especially for West Germany.
Notably, for Austria, the penalty parameter is nearly zero, rendering the penalized SCM nearly identical to the original SCM. Overall, iSCM and penalized iSCM yield very similar results.
(a) Gaps West Germany
(b) Gaps Austria
(c) Penalized iSCM vs iSCM West Germany
(d) Penalized iSCM vs iSCM Austria
Figure 2: Estimated effects with Penalized SCM